Given: Q = 20K0.3 L0.7, C* = $1,000, r = $30 and w = $10. Determine the amount of labor and capital that maximizes output and determine the optimal amount of output. Calculate and interpret the Lagrangian multiplier.


1
Expert's answer
2021-06-01T14:02:24-0400
"30K+10L=1000"

"Q=20K^{0.3}L^{0.7}+\lambda (1000-30K-10L)"

"\frac{\delta Q}{\delta K}=6(\frac{L}{K})^{0.7}-30 \lambda"

"\frac {\delta Q}{\delta L}=14(\frac{K}{L})^{0.3}-10\lambda"

"L=7K"

"K=10"

"L=70"

"\lambda=0.78"